# Dark Photon Dark Matter and Low-Frequency Gravitational Wave Detection with Gaia-like Astrometry

### v1

@@ Appendix B Gauge invariance of the angular deflection
$h_{\mu\nu}(x)\rightarrow h_{\mu\nu}(x)-(\partial_{\mu}\xi_{\nu}+\partial_{\nu}\xi_{\mu})\;.$ (57)
@@
$=k^{\nu}_{(0)}\partial_{\nu}\xi^{\mu}(0)=\omega_{0}\mathcal{N}^{\nu}\partial_{\nu}\xi^{\mu}(0)\;.$ (58)
@@
From Eq. (49) we find
$\partial_{0}(\Delta\delta e^{\nu}_{i})=-\Delta\Gamma_{i0}^{\nu}=\partial_{0}\partial_{i}\xi^{\nu}\;,$ (59)
@@
$\Delta\delta n_{i}=-P_{ij}\left(\mathcal{N}^{\nu}\partial_{\nu}\xi_{j}-(\partial_{j}\xi_{\nu}+\partial_{\nu}\xi_{j})\mathcal{N}^{\nu}+\mathcal{N}_{\mu}\partial_{j}\xi^{\mu}\right)=0\;.$ (60)
@@
$\partial_{0}\delta e^{\nu}_{\mu}=-\Gamma_{\mu 0}^{\nu}=-\frac{1}{2}\partial_{0}h_{\mu}^{\nu}\;,$ (61)
@@
$=-P_{ij}\left(\frac{\omega_{0}}{\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\left(-n^{l}\partial_{0}h_{jl}+n^{l}n^{k}\partial_{k}h_{jl}-\frac{1}{2}n^{l}n^{k}\partial_{j}h_{kl}\right)\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}-\frac{1}{2}h_{jk}(0)n^{k}\right)\;.$ (63)
@@
$\frac{\mathrm{d}}{\mathrm{d}\lambda}h_{jl}=\omega_{0}\left(\partial_{0}h_{jl}-n^{k}\partial_{k}h_{jl}\right)\;,$ (64)
we can integrate the first two terms in Eq. (63)
@@
$=-P_{ij}\left(-\frac{1}{\lambda_{s}}\int_{0}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}-\frac{\omega_{0}}{2\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}n^{l}n^{k}\partial_{j}h_{kl}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}+\frac{1}{2}h_{jk}(0)n^{k}\right)\;.$ (65)
@@
$\delta n_{1i}$ $=\frac{1}{\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}\;,$ (66)
$\delta n_{2i}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}\;,$ (67)
$\delta n_{3i}$ $=-\frac{1}{2}P_{ij}h_{jk}(0)n^{k}\;.$ (68)
Now we estimate the relative size of these three terms. We introduce the coherent length $l_{c}$ which is $l_{c}=2\pi/(m_{A}v_{\rm vir})$. Without loss of generality, we consider the metric perturbation that takes the following form
$h_{ij}(x,t)=c\mathcal{H}_{ij}\rho(\boldsymbol{x})\cos(2m_{A}t+2\alpha)\;.$ (69)
where $\mathcal{H}_{ij}$ is an order one tensor, $\rho(\boldsymbol{x})$ refers to the local dark matter energy density, $c=-8\pi G/m_{A}^{2}$. For ultralight dark matter, a core structure can form at the center of the Galaxy with a radius of $\sim l_{c}$. The energy density of the core can be much larger than the energy density outside the core. We expect that the photon emitted by a star inside the core may have larger $\delta n_{1i}$ and $\delta n_{2i}$ since part of the worldline of the photon is inside the core. However, we will see in the following calculation that even the star is inside the core, $\delta n_{3i}$ still dominates $\delta n_{i}$.
Consider the case that the source of photon is inside the core. We suppose that at $\lambda=\lambda_{b}$, the photon touches the boundary of the core. Therefore we expect $\omega_{0}|\lambda_{s}-\lambda_{b}|\sim l_{c}$. First we calculate $\delta n_{1i}$. We separate $\delta n_{1i}$ into two parts
$\delta n_{1i}=\frac{c}{\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}+\frac{c}{\lambda_{s}}P_{ij}\int_{0}^{\lambda_{b}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}\;,$ (70)
Since the $h_{ij}$ inside the core is much larger, we only calculate the first integration. In addition, the integral length is approximately one coherent length, so we can use Eq. (69) to express $h_{ij}$. Insert Eq. (69) into $\delta n_{1i}$, we obtain
$\delta n_{1i}=\frac{c}{\lambda_{s}}P_{ij}n^{l}\mathcal{H}_{jl}\int_{\lambda_{b}}^{\lambda_{s}}\rho(\lambda^{\prime})\cos(2m_{A}\omega_{0}\lambda^{\prime}+2\alpha)\mathrm{d}\lambda^{\prime}\sim P_{ij}n^{l}\mathcal{H}_{jl}\frac{c\rho_{c}}{m_{A}\omega_{0}\lambda_{s}}\;,$ (71)
where $\rho_{c}$ refers to the energy density of the soliton core. For simplicity, we use NFW profile to estimate $\rho_{c}$. Suppose the energy density of DPDM near the Earth is $\rho_{0}$. The NFW profile takes the form as McMillan (2011)
$\rho_{h}=\frac{\rho_{h,0}}{x(1+x)^{2}}\;,$ (72)
@@
$\frac{\rho_{c}}{\rho_{0}}\sim\frac{l_{e}}{l_{c}}\sim l_{e}m_{A}v_{\rm vir}\;,$ (73)
where $l_{e}=8\mathrm{kpc}$ is the distance between the Earth and the center of the Milky Way. Using the expression of $\rho_{c}$, we obtain
$\delta n_{1i}\sim P_{ij}n^{l}\mathcal{H}_{jl}\frac{c\rho_{0}v_{\rm vir}l_{e}}{x_{s}}\sim cP_{ij}n^{l}\mathcal{H}_{jl}v_{\rm vir}\rho_{0}\sim v_{\rm vir}\delta n_{3i}\;,$ (74)
$x_{s}$ is the distance from the star to the observer. $l_{e}/x_{s}$ is an order one parameter, so we ignore it. The virial velocity is $v_{\rm vir}\sim 10^{-3}$. Therefore, $\delta n_{1i}$ is suppressed compared to $\delta n_{3i}$.
@@
$\delta n_{2i}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}\int_{\lambda^{\prime\prime}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime}n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})$
$=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})$
$=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})+\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{b}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})\;.$ (75)
@@
Again we only calculate the first integral. Inside the core, we find $\omega_{0}|\lambda_{s}-\lambda^{\prime\prime}|\sim l_{c}=\frac{1}{m_{A}v_{\rm vir}}$ and $\partial_{i}h_{jk}\sim m_{A}v_{\rm vir}h_{jk}$. Then we have
$\delta n_{2i}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})$
$\sim\frac{1}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}\left(\frac{1}{m_{A}v_{\rm vir}}\right)n^{l}n^{k}m_{A}vh_{kl}(\lambda^{\prime\prime})$
$\sim\frac{1}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}n^{l}n^{k}h_{kl}(\lambda^{\prime\prime})\sim\delta n_{1i}\;,$ (76)
@@
where we find $\omega_{0}(\lambda_{s}-\lambda^{\prime\prime})$ and the spatial derivative of $h_{ij}$ cancel with each other, making $\delta n_{1i}$ and $\delta n_{2i}$ at the same order. Then similar to $\delta n_{1i}$, $\delta n_{2i}$ is also suppressed compared to $\delta n_{3i}$.

### v2

@@ Appendix B Gauge invariance of the angular deflection
$h_{\mu\nu}(x)\rightarrow h_{\mu\nu}(x)-(\partial_{\mu}\xi_{\nu}(x)+\partial_{\nu}\xi_{\mu}(x))\;.$ (B18)
@@
$=k^{\nu}_{(0)}\partial_{\nu}\xi^{\mu}(0)=\omega_{0}\mathcal{N}^{\nu}\partial_{\nu}\xi^{\mu}(0)\;.$ (B19)
@@
From Eq. (B10) we find
$\partial_{0}(\Delta\delta e^{\nu}_{i})=-\Delta\Gamma_{i0}^{\nu}=\partial_{0}\partial_{i}\xi^{\nu}\;,$ (B20)
@@
$\Delta\delta n_{i}=-P_{ij}\left(\mathcal{N}^{\nu}\partial_{\nu}\xi_{j}-(\partial_{j}\xi_{\nu}+\partial_{\nu}\xi_{j})\mathcal{N}^{\nu}+\mathcal{N}_{\mu}\partial_{j}\xi^{\mu}\right)=0\;.$ (B21)
@@
$\partial_{0}\delta e^{\nu}_{\mu}=-\Gamma_{\mu 0}^{\nu}=-\frac{1}{2}\partial_{0}h_{\mu}^{\nu}\;,$ (B22)
@@
$=-P_{ij}\left(\frac{\omega_{0}}{\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}\left(-n^{l}\partial_{0}h_{jl}+n^{l}n^{k}\partial_{k}h_{jl}-\frac{1}{2}n^{l}n^{k}\partial_{j}h_{kl}\right)\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}-\frac{1}{2}h_{jk}(0)n^{k}\right)\;.$ (B24)
@@
$\frac{\mathrm{d}}{\mathrm{d}\lambda}h_{jl}=\omega_{0}\left(\partial_{0}h_{jl}-n^{k}\partial_{k}h_{jl}\right)\;,$ (B25)
we can integrate the first two terms in Eq. (B24)
@@
$=-P_{ij}\left(-\frac{1}{\lambda_{s}}\int_{0}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}-\frac{\omega_{0}}{2\lambda_{s}}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}n^{l}n^{k}\partial_{j}h_{kl}\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}+\frac{1}{2}h_{jk}(0)n^{k}\right)\;.$ (B26)
@@
$\delta n_{1i}$ $=\frac{1}{\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}\;,$ (B27)
$\delta n_{2i}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\int_{0}^{\lambda^{\prime}}n^{l}n^{k}\partial_{j}h_{kl}(\lambda^{\prime\prime})\mathrm{d}\lambda^{\prime\prime}\mathrm{d}\lambda^{\prime}\;,$ (B28)
$\delta n_{3i}$ $=-\frac{1}{2}P_{ij}h_{jk}(0)n^{k}\;.$ (B29)
Now we estimate the relative size of these three terms. Without loss of generality, we consider the metric perturbation that takes the following form
$h_{ij}^{\rm DP}(\boldsymbol{x},t)=c\mathcal{H}_{ij}^{\rm DP}\rho(\boldsymbol{x})\cos(2m_{A}t+2\alpha)\;,$ (B30)
where $\mathcal{H}_{ij}^{\rm DP}$ is an order one tensor, $\rho(\boldsymbol{x})$ refers to the local dark matter energy density, $c=-8\pi G/m_{A}^{2}$. For ultralight dark matter, a core structure can form at the center of the Galaxy with a radius of $\sim l_{c}=\lambda_{dB}/(2\pi)=1/(m_{A}v_{\rm vir})$. The energy density of the core is a constant, which can be much larger than the energy density outside the core. We expect that the photon emitted by a star inside the core may have larger $\delta n_{1i}$ and $\delta n_{2i}$ since part of the worldline of the photon is inside the core. However, we will see in the following calculation that even when the star is inside the core, $\delta n_{3i}$ is still the dominant component of the total angular deflection $\delta n_{i}$.
B.1 Earth outside the core, star inside the core
Consider the case that the star is inside the core. It is valid for $l_{s}<\lambda_{\rm dB}=2\pi l_{c}<l_{e}$ (with $l_{e}=8\,{\rm kpc}$ and $l_{s}$ being respectively the Earth and star’s distance to the galactic center), implying $m_{A}\geq 10^{-23.4}\,{\rm eV}$. We suppose that at $\lambda=\lambda_{b}$, the photon touches the boundary of the core. Therefore we expect $\omega_{0}|\lambda_{s}-\lambda_{b}|\sim l_{c}$. First we calculate $\delta n_{1i}$. We separate $\delta n_{1i}$ into two parts
$\delta n_{1i}=\frac{c}{\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}+\frac{c}{\lambda_{s}}P_{ij}\int_{0}^{\lambda_{b}}n^{l}h_{jl}(\lambda^{\prime})\mathrm{d}\lambda^{\prime}\;.$ (B31)
Since the $h_{ij}$ inside the core is significantly larger, we only calculate the first integration. In addition, the integral length is approximately one coherent length, so we can use Eq. (B30) to express $h_{ij}$. Insert Eq. (B30) into $\delta n_{1i}$, we obtain
$\delta n_{1i}^{\rm DP}=\frac{c}{\lambda_{s}}P_{ij}n^{l}\mathcal{H}_{jl}^{\rm DP}\int_{\lambda_{b}}^{\lambda_{s}}\rho(\lambda^{\prime})\cos(2m_{A}\omega_{0}\lambda^{\prime}+2\alpha)\mathrm{d}\lambda^{\prime}\sim P_{ij}n^{l}\mathcal{H}_{jl}^{\rm DP}\frac{c\rho_{c}}{m_{A}\omega_{0}\lambda_{s}}\;,$ (B32)
where $\rho_{c}$ refers to the energy density of the soliton core. For simplicity, we use the NFW profile to estimate $\rho_{c}$. Suppose the energy density of DPDM near the Earth is $\rho_{0}$. The NFW profile takes the form as (McMillan (2011))
$\rho_{h}=\frac{\rho_{h,0}}{x(1+x)^{2}}\;,$ (B33)
@@ B.1 Earth outside the core, star inside the core
$\frac{\rho_{c}}{\rho_{0}}\sim\frac{l_{e}}{l_{c}}\sim l_{e}m_{A}v_{\rm vir}\;.$ (B34)
Using the expression of $\rho_{c}$, we obtain
$\delta n_{1i}^{\rm DP}\sim P_{ij}n^{l}\mathcal{H}_{jl}^{\rm DP}\frac{c\rho_{0}v_{\rm vir}l_{e}}{x_{s}}\sim cP_{ij}n^{l}\mathcal{H}_{jl}^{\rm DP}v_{\rm vir}\rho_{0}\sim v_{\rm vir}\delta n_{3i}^{\rm DP}\;,$ (B35)
$x_{s}$ is the distance from the star to the observer, $\delta n_{3i}^{\rm DP}=-\frac{1}{2}P_{ij}h_{jk}^{\rm DP}(0)n^{k}$. $l_{e}/x_{s}$ is an order one parameter, so we ignore it. The virial velocity is $v_{\rm vir}\sim 10^{-3}$. Therefore, $\delta n_{1i}$ is suppressed compared to $\delta n_{3i}$.
@@
$\delta n_{2i}^{\rm DP}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}\int_{\lambda^{\prime\prime}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime}n^{l}n^{k}\partial_{j}h_{kl}^{\rm DP}(\lambda^{\prime\prime})$
$=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}^{\rm DP}(\lambda^{\prime\prime})$
$=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}^{\rm DP}(\lambda^{\prime\prime})+\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{0}^{\lambda_{b}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}^{\rm DP}(\lambda^{\prime\prime})\;.$ (B36)
@@
Again we only calculate the first integral. Inside the core, we find $\omega_{0}|\lambda_{s}-\lambda^{\prime\prime}|\sim l_{c}=\frac{1}{m_{A}v_{\rm vir}}$ and $\partial_{i}h_{jk}^{\rm DP}\sim m_{A}v_{\rm vir}h_{jk}^{\rm DP}$. Then we have
$\delta n_{2i}^{\rm DP}$ $=\frac{\omega_{0}}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}(\lambda_{s}-\lambda^{\prime\prime})n^{l}n^{k}\partial_{j}h_{kl}^{\rm DP}(\lambda^{\prime\prime})$
$\sim\frac{1}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}\left(\frac{1}{m_{A}v_{\rm vir}}\right)n^{l}n^{k}m_{A}v_{\rm vir}h_{kl}^{\rm DP}(\lambda^{\prime\prime})$
$\sim\frac{1}{2\lambda_{s}}P_{ij}\int_{\lambda_{b}}^{\lambda_{s}}\mathrm{d}\lambda^{\prime\prime}n^{l}n^{k}h_{kl}^{\rm DP}(\lambda^{\prime\prime})\sim\delta n_{1i}^{\rm DP}\;,$ (B37)
@@
where we find $\omega_{0}(\lambda_{s}-\lambda^{\prime\prime})$ and the spatial derivative of $h_{ij}^{\rm DP}$ cancel with each other, making $\delta n_{1i}^{\rm DP}$ and $\delta n_{2i}^{\rm DP}$ at the same order. Then similar to $\delta n_{1i}^{\rm DP}$, $\delta n_{2i}^{\rm DP}$ is also suppressed compared to $\delta n_{3i}^{\rm DP}$.
B.2 Both the Earth and the star are inside the core
We assume that the star and the Earth are inside the core. It requires $m_{A}<10^{-23.4}$ to include the Earth inside the core. We treat the dark matter density at the core as a constant $\rho(\boldsymbol{x})=\rho_{c}$. Then the first term $\delta n_{1i}$ in Eq. (B29) has the approximated value
$\frac{\delta n_{1i}^{\rm DP}}{\delta n_{3i}^{\rm DP}}$ $\leq\frac{l_{\rm Comp}}{D}\sin\left(\frac{D}{l_{\rm Comp}}\right)\sim\begin{cases}l_{\rm Comp}/D\,,&l_{\rm Comp}\ll D\\ 1\,,&l_{\rm Comp}\gg D\end{cases}\,,$ (B38)
where $\omega_{0}\lambda_{s}=D$ is the distance of the star, $1/m_{A}=l_{\rm Comp}$ is the Compton wavelength of the dark photon field. We find that $\delta n_{1i}$ consistently smaller than $\delta n_{3i}$ when $l_{\rm Comp}\ll D$, which holds true for $m_{A}\geq 10^{-24}\,{\rm eV}$ with the typical star distance of $1\,{\rm kpc}$. For $\delta n_{2i}^{\rm DP}$, we also easily find that it is at the same order of $\delta n_{2i}^{\rm DP}$ for the same reason we described in the previous sub-section.
B.3 Both the Earth and the star are outside the core
At last, we consider when the star and the Earth are outside the core. It is valid for a larger dark matter mass $m_{A}\gtrsim 10^{-23}\,{\rm eV}$, depending on the location of the star. Without loss of generality, we assume that the star lies on the line between the Earth and the galaxy center and the boundary of the core. Then, from previous sections, we have the dark matter density at the star $\rho_{s}=\rho_{c}\simeq\rho_{0}l_{c}/l_{e}$ and
$\frac{\delta n_{1i}^{\rm DP}}{\delta n_{3i}^{\rm DP}}\lesssim\frac{l_{e}}{l_{c}}\frac{l_{\rm Comp}}{D}\sin\left(\frac{D}{l_{\rm Comp}}\right)\sim\begin{cases}v_{\rm vir}\l_{e}/D\,,&l_{\rm Comp}\ll D\\ l_{e}/l_{c}\,\,,&l_{\rm Comp}\gg D\end{cases}\,.$ (B39)
where we used $D\gg l_{\rm Comp}$, $l_{c}=v_{\rm vir}l_{\rm Comp}$ and $l_{e}\simeq 8\,{\rm kpc}$ and $D\sim 1\,{\rm kpc}$. Similarly, $\delta n_{2i}^{\rm DP}$ is at the same order of $\delta n_{1i}^{\rm DP}$. As we estimated before, with the mass $m_{A}\geq 10^{-24}\,{\rm eV}$, we always have $l_{\rm Comp}\ll D$. Considering that $v_{\rm vir}\simeq 10^{-3}$, it is safe to ignore $\delta n_{1i}$ and $\delta n_{2i}$.